Tim is 6 years older than JoAnn. Seven years from now the sum of their ages will be 78.
Answer by richwmiller(17219)

tim=jo+6
tim+7+jo+7=78
jo+6+7+jo+7=78
jo=(78-(14)-6)/2
jo=(58)/2
jo=29
tim=29+6
tim=35
jo=29 tim= 35